# Magnetic disk winds in protoplanetary disks:

### v2

@@ 2.2 Magnetic disk wind model
${{\rm C}_{{\rm FUV},T_{z\phi}}}=\frac{1}{2}\bigg{(}\frac{L_{T}}{L_{\odot}}\bigg{)}^{0.5},$ (20)
in the wind mass loss rate and surface stress equations, respectively. Here $L_{T}$ is total, i.e., accretion plus stellar luminosity of the system. The exponent for the ${\rm C}_{{\rm FUV},\dot{M}_{w}}$ term is chosen to be marginally lower than unity, since a larger value results in a significant decrease in the wind velocity during an outburst (see Appendix A). We note that this dependence incorporates several assumptions implicitly, e.g., the FUV penetration depth is proportional to the central FUV luminosity, which in turn is proportional to the total luminosity of the accreting star, along with the exact exponents of the wind mass loss and stress. Additionally, the dusty winds can self-shield the FUV irradiation; the winds from the inner regions can absorb FUV photons and interfere with the wind launching from the remainder of the disk (Panoglou et al., 2012; Rab et al., 2022). With the exponents in the above equations being less than unity, self-shielding is inherently assumed in the model. We neglect any additional uncertainties, e.g., effects of photochemistry in the upper layers on the ionization (Bergin et al., 2003; Woitke et al., 2009). Inclusion of these terms (Eqs. (19) and (20)) results in an increased wind mass loss rate during outbursts that is approximately an order of magnitude less than the accretion rate onto the star. A caveat of this approach is that the radiation takes a finite time to reach the disk, which should cause a delay in producing magnetic winds. However, this delay is a small fraction of the dynamical time at a given radius and thus, can be safely neglected.
As the winds originate centrifugally from a protoplanetary disk, it is essential to determine the precise extent of this disk. During the collapse of molecular cloud to form a star, a gas parcel of specific angular momentum $j$ will be accelerated towards the central star in a parabolic orbit. This orbit intersects the plane of the disk at
@@
where $M_{\rm enc}$ is the mass enclosed within the orbit of the gas parcel (Dominik, 2015). A gas parcel can be part of the centrifugal disk only if it has sufficient specific angular momentum to resist the inward fall. This criterion implies that material at a given radius $r$ forms a centrifugal disk, if the ratio $r/r_{\rm eq}$ is less than unity. In a protoplanetary disk, however, the gas experiences additional pressure support and hence, it moves with a sub-Keplerian velocity. Thus, for calculating the outer boundary of the centrifugal disk, we use a less stringent criterion of $r/r_{\rm eq}<1.2$. The disk winds are active only within this region and are completely inactive outside. In addition, the disk winds are turned off below a gas surface density threshold of $0.1{\rm g\,cm^{-2}}$, in order to limit unintended consequences of the winds at large distances. We note that due to the diminishing gas surface density, the wind mass loss rate and the surface stress decrease with radius and become progressively insignificant. We obtain a single “centrifugal radius” ($R_{\rm cf}$) by azimuthally averaging the extent of the centrifugal disk at a given time.
The action of the disk winds needs to be taken into account self-consistently in the hydrodynamic equations that are solved in FEOSAD. For this, the equations that need to be modified are - mass continuity equation for gas and small dust (Eqs. (1) and (4)), momentum equation (Eq. (2.1)) and energy transport equation (Eq. (3)). For solving the MHD equations, FEOSAD uses operator splitting technique, which divides problem into two partial substeps- the source and the transport step (e.g., Stone & Norman, 1992). During the numerical calculations, the corrections due to wind mass loss and surface stress ($\dot{\Sigma}_{w}$ and $T_{z\phi}^{\rm Max}$) are computed in each cell at the beginning of the source step. Corrections to gas and small dust surface density as well as energy are done at the end of the source step, but before the transport step begins. A fraction, $d\Sigma_{w,g}$, is removed from the gas surface density, such that
@@
In order to prevent sudden, large mass loss during a timestep, the mass lost due to the wind is capped at 5% of the gas mass contained within the cell. In such a case, the disk $T_{z\phi}^{\rm Max}$ is capped such that the velocity of the escaping wind remains unchanged (See Appendix A for details).
@@
where $C_{\dot{M}}=1.54\times 10^{-5}$ and $C_{T_{z\phi}}=5.8\times 10^{-5}$ are obtained by consolidating all other constants in the respective equations. With this framework of magnetic disk winds, all of the exponents in Eqs. (25) and (26) are informed by the shearing box simulations. The only “free” parameters are $C_{H}$ and $C_{\beta}$, which arise from the discrepancy with respect to non-ideal MHD effects and uncertainties in the wind mass loss rate, respectively. We will constrain these parameters with the help of observations in Section 3.3. We note that with the assumption of spatially uniform mass to flux ratio, $\lambda=\Sigma_{\rm g}/B_{\rm z}$, $\dot{\Sigma}_{w,g}\sim\Sigma_{\rm g}c_{\rm s}$ and $T_{z\phi}^{\rm Max}\sim\Sigma_{\rm g}^{5/4}c_{\rm s}^{2/3}$, which explains how the disk wind properties vary with the local quantities.
@@ 2.3 Interface to ProDiMo
The main challenge for this interface is that FEOSAD uses 2D polar coordinates $(r,\phi)$, considering vertically integrated quantities, whereas ProDiMo uses 2D cylinder coordinates $(r,z)$, assuming the disk to be axisymmetric. We have therefore developed the following scheme. First, the FEOSAD disk structure at a given time is azimuthally averaged. A PPD structure is not strictly axisymmetric, however, the azimuthal asymmetries get increasingly smoothed out as the disk evolves and the GI diminishes. Second, the following quantities are passed to ProDiMo: the gas column density $\Sigma_{\rm g}$, the column density of small dust $\Sigma_{\rm d,sm}$, the column density of grown dust $\Sigma_{\rm d,gr}$, the viscosity parameter $\alpha_{\rm turb}$, the scale height $H_{\rm g}$, and the maximum grain radius $a_{\rm max}$. All these quantities are passed to ProDiMo as vectors on approximately 200 log-equidistant radial grid points used by FEOSAD. In addition, the stellar luminosity $L_{\star}$, the effective temperature of the star $T_{\rm eff}$ and the stellar mass $M_{\star}$ are passed as scalars. Third, ProDiMo sets up the 2D structure of the gas disk in $(r,z)$ on the provided radial grid points, while adding a few additional grid points by extrapolation near the inner rim, so that the penetration of stellar light into the disk is traced correctly. We then use the radius-dependent scale height $H_{g}(r)$ to set up the vertical gas density structure $\rho(r,z)$ at each radius using a Gaussian function.
@@
The dust grains in FEOSAD have size distribution between a fixed minimum size $a_{\rm min}\!=\!0.005\,\mu$m and a variable maximum size $a_{\rm max}$, while assuming a continuous power-law size distribution of index –3.5. In each vertical column, ProDiMo sets up the unsettled grain size distribution function using 100 size bins, and then the grains are settled according to the prescription of Riols & Lesur (2018) in each bin with the settling parameter $\alpha_{\rm turb}$ (Woitke et al., 2024). We use the DIANA standard dust opacities Woitke et al. (2016), assuming the grains to be composed of 60% silicate, 15% amorphous carbon, with 25% porosity and without any ice. It is possible that ice formation can locally increase the dust opacities by more than a factor 100, especially at UV to mid-IR wavelengths (Arabhavi et al., 2022). However, the ice exists only deep in the disk ($A_{V}\!\gtrsim\!10$) where it is protected from UV photons, i.e., where the refractory dust is optically thick. Therefore, ice opacities do not have a strong influence on the resulting dust temperature structure and the spectral appearance of the disk concerning CO sizes and millimeter fluxes. Although some uncertainties remain when vertical mixing is included in the models (Woitke et al., 2022), in which case the grains can be mixed up faster than photodesorption can destroy the ice. Based on the settled dust structure, stellar and interstellar irradiation, and internal viscous heating, ProDiMo performs continuum radiative transfer calculations. This results in the disk internal dust temperature structure $T_{\rm d}(r,z)$ and produces ray-traced continuum images and CO line maps using a polar grid in the image plane with $225\times 72$ segments.
@@
The viscosity parameter $\alpha_{\rm turb}(r)$ is also used for computing the viscous heating in ProDiMo. Here we use a 2D diffusion solver in the optically thick core of the disk as described in the Appendix of Oberg et al. (2022), and use the following formula for the local dust heating (Frank et al., 2002)
@@ 2.4 Initial conditions and model parameters
The MHD simulations in FEOSAD start with the gravitational collapse phase of a starless molecular cloud core. The initial mass distribution as well as angular velocity is consistent with axisymmetric core collapse, where the angular momentum remains constant and magnetic fields are expelled due to ambipolar diffusion. The collapse of the core continues such that a star is formed within the central sink of the computational domain. A surrounding protostellar disk simultaneously forms while the envelope is infalling and the system continues to evolve through the embedded phase. For a detailed description of the initial structure and angular velocity profile, see Basu (1997) and Vorobyov et al. (2020b). Since the focus of this study is on the effects of the magnetic winds on the disk formation and its evolution, all simulations start with identical initial conditions, which are summarized in Table 1. The cloud core of 0.83 $M_{\odot}$ forms the initial gas mass reservoir and results in a young star of mass between $0.5-0.7M_{\odot}$, depending on the efficiency of the particular model. The ratio of kinetic to gravitational energy is set to $E_{\rm rot}/E_{\rm grav}=0.23\%$, consistent with the observations of prestellar cores (Caselli et al., 2002; Li et al., 2012). The disks formed with these conditions are gravitationally unstable, while being stable to fragmentation and clump formation (Vorobyov, 2013). These particular initial conditions are chosen because formation of disks in similar conditions is extensively studied with FEOSAD simulations (e.g., Kadam et al., 2020; Vorobyov et al., 2020b). Additionally, we will compare our results with ALMA survey of low mass star forming regions, wherein the stellar mass of the sample centers at M0 type star of mass $\approx 0.5$ $M_{\odot}$ (Ansdell et al., 2016; Manara et al., 2023a).
@@
The initial bulk dust-to-gas mass ratio, $\zeta_{\rm d2g}$, is set to the canonical interstellar medium (ISM) value of 1% and all the dust mass is in the fully coupled submicron-sized small particles. The initial cloud core gas temperature is 20 K, which was also the temperature of the background radiation (Kennicutt & Evans, 2012). A uniform background magnetic field of $10^{-5}$ G normal to the plane of the disk is also assumed (Liu et al., 2021). The mass-to-flux ratio ($\lambda_{c}=\Sigma_{g}/B_{z}$, in the units of the critical value $1/2\pi G^{1/2}$) in dense star-forming cores is observed to be $\lambda_{c}\simeq 2$ (Crutcher, 2012), with a cloud or core being supercritical for values of $\lambda_{c}$ above unity (Mouschovias & Spitzer, 1976; Nakano & Nakamura, 1978). A commensurate value of $\lambda_{c}=2$ was used for our earlier investigations, however, some flux should be lost during the collapse phase due to ambipolar diffusion (Shu et al., 1987; Dapp et al., 2012). Since FEOSAD does not include non-ideal MHD effects, we can justify a somewhat larger value of $\lambda_{c}=5$, and thus, partially compensate for the model limitations. The resolution of all simulations is $184\times 128$ in polar coordinates ($r,\phi$), with a logarithmic spacing in the radial direction and uniform spacing in the azimuthal direction at each radius. The increased resolution in the $r$-direction is chosen to minimize the elongation of grid cells, preventing potential numerical artifacts. The convergence of FEOSAD code at different spatial resolutions has been previously confirmed (e.g. Kadam et al., 2021) and the current resolution sets $dr=2.6\times 10^{-2}$ au at the inner computational boundary at 0.53 au. The location of this inner boundary or sink cell excludes the actual inner disk radius of a PPD occurring at $\approx 0.1$ au, which is determined by the magnetospheric accretion or dust sublimation radius (Dullemond & Monnier, 2010; Hartmann et al., 2016). However, the marginally larger boundary relaxes the Courant condition for computational feasibility of the simulations (Courant et al., 1928), while being sufficient for capturing important dynamical phenomena occurring in the inner disk, such as episodic accretion. A carefully constructed inflow-outflow boundary condition is implemented at the inner sink cell in order to prevent the artificial drop in the gas surface density (Vorobyov et al., 2018a; Kadam et al., 2022).
Table 1: Initial parameters common across simulations
@@
Table 2: List of simulations
@@
In this study, we present results from six simulations as listed in Table 2. The fiducial simulations, GV-2 and GV-4, evolve via the combined action of gravitational and viscous torques. The wind quantities in these models are calculated, however, no corrections are made to the disk evolution equations. The value of background $\alpha_{\rm MRI}$ in GV-2 is 0.01. As discussed in Sect. 1, protoplanetary disks may exhibit significantly lower value of turbulent $\alpha$ and this possibility is explored in model GV-4, with $\alpha_{\rm MRI}=10^{-4}$. Model WI-2 is identical to GV-2, wherein the disk winds are taken into account self-consistently. The gravitoviscous forces are active in all disk models, including those where the magnetic winds are included. In the case of WI-4, winds are considered with a lower value of $\alpha_{\rm MRI}$ and thus, it is otherwise identical with GV-4. Model WI-3 considers winds with the intermediate value of $\alpha_{\rm MRI}=10^{-3}$. In the case of WI-3a, the wind parameters $C_{H}$ and $C_{\beta}$ are adjusted, in order to demonstrate how a better fit to the observations can be obtained. With this limited parameter space study some of the most important possibilities in the disk wind scenario are explored.

### v3

@@ 2.2 Magnetic disk wind model
${{\rm C}_{{\rm FUV},T_{z\phi}}}=\frac{1}{2}\bigg{(}\frac{L_{T}}{L_{\odot}}\bigg{)}^{0.5}.$ (20)
Here, $L_{T}$ is total, that is, the accretion plus stellar luminosity of the system. The exponent for the ${\rm C}_{{\rm FUV},\dot{M}_{w}}$ term is chosen to be marginally lower than unity, since a larger value results in a significant decrease in the wind velocity during an outburst (see Appendix A). We note that this dependence incorporates several assumptions implicitly, for example, the FUV penetration depth is proportional to the central FUV luminosity, which in turn is proportional to the total luminosity of the accreting star, along with the exact exponents of the wind mass loss and stress. Additionally, the dusty winds can self-shield the FUV irradiation; the winds from the inner regions can absorb FUV photons and interfere with the wind launching from the remainder of the disk (Panoglou et al., 2012; Rab et al., 2022). With the exponents in the above equations being less than unity, self-shielding is inherently assumed in the model. We neglect any additional uncertainties, for example, effects of photochemistry in the upper layers on the ionization (Bergin et al., 2003; Woitke et al., 2009). Inclusion of these terms (Eqs. (19) and (20)) results in an increased wind mass loss rate during outbursts that is approximately an order of magnitude less than the accretion rate onto the star. A caveat of this approach is that the radiation takes a finite time to reach the disk, which should cause a delay in producing magnetic winds. However, this delay is a small fraction of the dynamical time at a given radius, and thus, it can be safely neglected.
As the winds originate centrifugally from a protoplanetary disk, it is essential to determine the precise extent of this disk. During the collapse of molecular cloud to form a star, a gas parcel of specific angular momentum $j$ will be accelerated toward the central star in a parabolic orbit. This orbit intersects the plane of the disk at
@@
where $M_{\rm enc}$ is the mass enclosed within the orbit of the gas parcel (Dominik, 2015). A gas parcel can be part of the centrifugal disk only if it has sufficient specific angular momentum to resist the inward fall. This criterion implies that material at a given radius $r$ forms a centrifugal disk, if the ratio $r/r_{\rm eq}$ is less than unity. In a protoplanetary disk, however, the gas experiences additional pressure support, and hence, it moves with a sub-Keplerian velocity. Thus, for calculating the outer boundary of the centrifugal disk, we use a less stringent criterion of $r/r_{\rm eq}<1.2$. The disk winds are active only within this region and are completely inactive outside. In addition, the disk winds are turned off below a gas surface density threshold of $0.1{\rm g\,cm^{-2}}$, in order to limit unintended consequences of the winds at large distances. We note that due to the diminishing gas surface density, the wind mass loss rate and the surface stress decrease with radius and become progressively insignificant. We obtain a single “centrifugal radius” ($R_{\rm cf}$) by azimuthally averaging the extent of the centrifugal disk at a given time.
The action of the disk winds needs to be taken into account self-consistently in the hydrodynamic equations that are solved in FEOSAD. For this, the equations that need to be modified are - mass continuity equation for gas and small dust (Eqs. (1) and (4)), momentum equation (Eq. (2.1)) and energy transport equation (Eq. (3)). For solving the MHD equations, FEOSAD uses operator splitting technique, which divides problem into two partial substeps- the source and the transport step (e.g., Stone & Norman, 1992). During the numerical calculations, the corrections due to wind mass loss and surface stress ($\dot{\Sigma}_{w}$ and $T_{z\phi}^{\rm Max}$) are computed in each cell at the beginning of the source step. Corrections to gas and small dust surface density as well as energy are done at the end of the source step, but before the transport step begins. A fraction, $d\Sigma_{w,g}$, is removed from the gas surface density such that
@@
In order to prevent a sudden, large amount of mass loss during a time step, the mass lost due to the wind is capped at 5% of the gas mass contained within the cell. In such a case, the disk $T_{z\phi}^{\rm Max}$ is capped such that the velocity of the escaping wind remains unchanged (see Appendix A for details).
@@
where $C_{\dot{M}}=1.54\times 10^{-5}$ and $C_{T_{z\phi}}=5.8\times 10^{-5}$ are obtained by consolidating all other constants in the respective equations. With this framework of magnetic disk winds, all of the exponents in Eqs. (25) and (26) are informed by the shearing box simulations. The only “free” parameters are $C_{H}$ and $C_{\beta}$, which arise from the discrepancy with respect to non-ideal MHD effects and uncertainties in the wind mass loss rate, respectively. We constrain these parameters with the help of observations in Sect. 3.3. We note that with the assumption of spatially uniform mass to flux ratio, $\lambda=\Sigma_{\rm g}/B_{\rm z}$, $\dot{\Sigma}_{w,g}\sim\Sigma_{\rm g}c_{\rm s}$ and $T_{z\phi}^{\rm Max}\sim\Sigma_{\rm g}^{5/4}c_{\rm s}^{2/3}$, which explains how the disk wind properties vary with the local quantities.
@@ 2.3 Interface to ProDiMo
The main challenge for this interface is that FEOSAD uses 2D polar coordinates $(r,\phi)$, considering vertically integrated quantities, whereas ProDiMo uses 2D cylinder coordinates $(r,z)$, assuming the disk to be axisymmetric. We have therefore developed the following scheme. First, the FEOSAD disk structure at a given time is azimuthally averaged. A PPD structure is not strictly axisymmetric; however, the azimuthal asymmetries increasingly become smoothed out as the disk evolves and the GI diminishes. Second, the following quantities are passed to ProDiMo: the gas column density $\Sigma_{\rm g}$, the column density of small dust $\Sigma_{\rm d,sm}$, the column density of grown dust $\Sigma_{\rm d,gr}$, the viscosity parameter $\alpha_{\rm turb}$, the scale height $H_{\rm g}$, and the maximum grain radius $a_{\rm max}$. All these quantities are passed to ProDiMo as vectors on approximately 200 log-equidistant radial grid points used by FEOSAD. In addition, the stellar luminosity $L_{\star}$, the effective temperature of the star $T_{\rm eff}$ and the stellar mass $M_{\star}$ are passed as scalars. Third, ProDiMo sets up the 2D structure of the gas disk in $(r,z)$ on the provided radial grid points, while adding a few additional grid points by extrapolation near the inner rim so that the penetration of stellar light into the disk is traced correctly. We then use the radius-dependent scale height $H_{g}(r)$ to set up the vertical gas density structure $\rho(r,z)$ at each radius using a Gaussian function.
@@
The dust grains in FEOSAD have size distribution between a fixed minimum size $a_{\rm min}\!=\!0.005\,\mu$m and a variable maximum size $a_{\rm max}$, while assuming a continuous power-law size distribution of index –3.5. In each vertical column, ProDiMo sets up the unsettled grain size distribution function using 100 size bins, and then the grains are settled according to the prescription of Riols & Lesur (2018) in each bin with the settling parameter $\alpha_{\rm turb}$ (Woitke et al., 2024). We use the DIANA standard dust opacities Woitke et al. (2016), assuming the grains to be composed of 60% silicate, 15% amorphous carbon, with 25% porosity and without any ice. It is possible that ice formation can locally increase the dust opacities by more than a factor of 100, especially at UV to mid-IR wavelengths (Arabhavi et al., 2022). However, the ice exists only deep in the disk ($A_{V}\!\gtrsim\!10$), where it is protected from UV photons, that is, where the refractory dust is optically thick. Therefore, ice opacities do not have a strong influence on the resulting dust temperature structure and the spectral appearance of the disk concerning CO sizes and millimeter fluxes. Although some uncertainties remain when vertical mixing is included in the models (Woitke et al., 2022), in which case the grains can be mixed up faster than photodesorption can destroy the ice. Based on the settled dust structure, stellar and interstellar irradiation, and internal viscous heating, ProDiMo performs continuum radiative transfer calculations. This results in the disk internal dust temperature structure $T_{\rm d}(r,z)$ and produces ray-traced continuum images and CO line maps using a polar grid in the image plane with $225\times 72$ segments.
@@
The viscosity parameter $\alpha_{\rm turb}(r)$ is also used for computing the viscous heating in ProDiMo. Here we use a 2D diffusion solver in the optically thick core of the disk as described in the appendix of Oberg et al. (2022) and use the following formula for the local dust heating (Frank et al., 2002),
@@ 2.4 Initial conditions and model parameters
The MHD simulations in FEOSAD start with the gravitational collapse phase of a starless molecular cloud core. The initial mass distribution as well as angular velocity is consistent with axisymmetric core collapse, where the angular momentum remains constant and magnetic fields are expelled due to ambipolar diffusion. The collapse of the core continues such that a star is formed within the central sink of the computational domain. A surrounding protostellar disk simultaneously forms while the envelope is infalling and the system continues to evolve through the embedded phase. Detailed description of the initial structure and angular velocity profile can be found in Basu (1997) and Vorobyov et al. (2020b). Since the focus of this study is on the effects of the magnetic winds on the disk formation and its evolution, all simulations start with identical initial conditions, which are summarized in Table 1. The cloud core of 0.83 $M_{\odot}$ forms the initial gas mass reservoir and results in a young star of mass between $0.5-0.7M_{\odot}$, depending on the efficiency of the particular model. The ratio of kinetic to gravitational energy is set to $E_{\rm rot}/E_{\rm grav}=0.23\%$, consistent with the observations of prestellar cores (Caselli et al., 2002; Li et al., 2012). The disks formed with these conditions are gravitationally unstable, while being stable to fragmentation and clump formation (Vorobyov, 2013). These particular initial conditions are chosen because formation of disks in similar conditions is extensively studied with FEOSAD simulations (e.g., Kadam et al., 2020; Vorobyov et al., 2020b). Additionally, we compare our results with the ALMA survey of low mass star forming regions, wherein the stellar mass of the sample centers at M0 type star of mass $\approx 0.5$ $M_{\odot}$ (Ansdell et al., 2016; Manara et al., 2023a).
@@
The initial bulk dust-to-gas mass ratio, $\zeta_{\rm d2g}$, is set to the canonical interstellar medium (ISM) value of 1% and all the dust mass is in the fully coupled submicron-sized small particles. The initial cloud core gas temperature is 20 K, which was also the temperature of the background radiation (Kennicutt & Evans, 2012). A uniform background magnetic field of $10^{-5}$ G normal to the plane of the disk is also assumed (Liu et al., 2021). The mass-to-flux ratio ($\lambda_{c}=\Sigma_{g}/B_{z}$, in the units of the critical value $1/2\pi G^{1/2}$) in dense star-forming cores is observed to be $\lambda_{c}\simeq 2$ (Crutcher, 2012), with a cloud or core being supercritical for values of $\lambda_{c}$ above unity (Mouschovias & Spitzer, 1976; Nakano & Nakamura, 1978). A commensurate value of $\lambda_{c}=2$ was used for our earlier investigations; however, some flux should be lost during the collapse phase due to ambipolar diffusion (Shu et al., 1987; Dapp et al., 2012). Since FEOSAD does not include non-ideal MHD effects, we can justify a somewhat larger value of $\lambda_{c}=5$ and thus partially compensate for the model limitations. The resolution of all simulations is $184\times 128$ in polar coordinates ($r,\phi$), with a logarithmic spacing in the radial direction and uniform spacing in the azimuthal direction at each radius. The increased resolution in the $r$-direction is chosen to minimize the elongation of grid cells, preventing potential numerical artifacts. The convergence of FEOSAD code at different spatial resolutions has been previously confirmed (e.g., Kadam et al., 2021) and the current resolution sets $dr=2.6\times 10^{-2}$ au at the inner computational boundary at 0.53 au. The location of this inner boundary or sink cell excludes the actual inner disk radius of a PPD occurring at $\approx 0.1$ au, which is determined by the magnetospheric accretion or dust sublimation radius (Dullemond & Monnier, 2010; Hartmann et al., 2016). However, the marginally larger boundary relaxes the Courant condition for computational feasibility of the simulations (Courant et al., 1928), while being sufficient for capturing important dynamical phenomena occurring in the inner disk, such as episodic accretion. A carefully constructed inflow-outflow boundary condition is implemented at the inner sink cell in order to prevent the artificial drop in the gas surface density (Vorobyov et al., 2018a; Kadam et al., 2022).
Table 1: Initial parameters common across simulations.
@@
Table 2: List of simulations.
@@
In this study, we present results from six simulations as listed in Table 2. The fiducial simulations, GV-2 and GV-4, evolve via the combined action of gravitational and viscous torques. The wind quantities in these models are calculated; however, no corrections are made to the disk evolution equations. The value of background $\alpha_{\rm MRI}$ in GV-2 is 0.01. As discussed in Sect. 1, protoplanetary disks may exhibit significantly lower value of turbulent $\alpha$ and this possibility is explored in model GV-4, with $\alpha_{\rm MRI}=10^{-4}$. Model WI-2 is identical to GV-2, wherein the disk winds are taken into account self-consistently. The gravitoviscous forces are active in all disk models, including those where the magnetic winds are included. In the case of WI-4, winds are considered with a lower value of $\alpha_{\rm MRI}$, and thus, it is otherwise identical with GV-4. Model WI-3 considers winds with the intermediate value of $\alpha_{\rm MRI}=10^{-3}$. In the case of WI-3a, the wind parameters $C_{H}$ and $C_{\beta}$ are adjusted, in order to demonstrate how a better fit to the observations can be obtained. With this limited parameter space study some of the most important possibilities in the disk wind scenario are explored.